Find to 3 decimal places the value of t for the t distribution for the following. Area in the left tail = 0.005 and n = 25 . t =
Answer #1

Sample size =n=25

Degree of freedom =n-1 = 25-1 = 24

o0.005

t critical value is -2.797(for left tail)